OpenStudy (anonymous):

implicit derivative of sin(2xy)

OpenStudy (anonymous):

I found the answer to be (-ycos(2xy))/(xcos(2xy))

OpenStudy (marissalovescats):

Your answer would have to be in terms of d/dy=

OpenStudy (marissalovescats):

Yes? That's the whole point of implicit differentiation.

OpenStudy (anonymous):

or y'

OpenStudy (anonymous):

Chain rule and product rule.

OpenStudy (marissalovescats):

No not y', y' would just be a regular derivative. This is implicit differentiation

OpenStudy (anonymous):

Yes I know, I'm telling you this problem requires chain rule and product rule though. Just with y'.

OpenStudy (marissalovescats):

Yes

OpenStudy (marissalovescats):

First thing you have to do is multiply the equation by d/dx

OpenStudy (marissalovescats):

@jim_thompson5910

OpenStudy (marissalovescats):

I believe u= 2xy Then you'd have cosu(dericative of u) Which you'd have to do implicit differentiation ad the product rule

jimthompson5910 (jim_thompson5910):

In the original problem, is sin(2xy) equal to anything?

OpenStudy (marissalovescats):

^ that's also why I was confused with this

OpenStudy (anonymous):

just f(x)

jimthompson5910 (jim_thompson5910):

so perhaps they meant to say y = sin(2xy)

jimthompson5910 (jim_thompson5910):

derive both sides to get y = sin(2xy) dy/dx = cos(2xy) * ( d/dx[ 2xy ] ) dy/dx = cos(2xy) * ( 2y + 2x*dy/dx ) and then isolate dy/dx

jimthompson5910 (jim_thompson5910):

Isolating dy/dx gives you... dy/dx = cos(2xy) * ( 2y + 2x*dy/dx ) dy/dx = 2y*cos(2xy) + cos(2xy)*2x*dy/dx dy/dx - cos(2xy)*2x*dy/dx = 2y*cos(2xy) dy/dx [1 - 2x*cos(2xy) ] = 2y*cos(2xy) dy/dx = ( 2y*cos(2xy) )/(1 - 2x*cos(2xy) ) so here's what it looks like in a cleaner format \[\Large \frac{dy}{dx} = \frac{2y\cos(2xy)}{1 - 2x\cos(2xy)}\]

jimthompson5910 (jim_thompson5910):

Optionally, you can think of dy/dx = cos(2xy) * ( 2y + 2x*dy/dx ) as z = q * ( 2y + 2x*z ) where z = dy/dx and q = cos(2xy) Then solve for z